a passenger train takes 2 hours 30 minutes more than an express train for a journey of 600km .the speed of the express train is 8km/h than the passenger train . Find the speed of the passenger train
1
Expert's answer
2019-06-12T02:28:11-0400

Let x be the speed of the passenger train,

then x+8 is the speed of the express train.

Time spent by passenger train was "600\/x" and time for the express train was 600/(x+8).

"600\/x -600\/(x+8)= 2.5"

"600\/x-600\/(x+8)-2.5=0"

"(600(x+8)-600x-2.5x(x+8))\/(x(x+8))=0"

"600x+4800-600x-2.5x^2-20x=0"

"-2.5x^2-20x+4800=0"

"x^2+8x-1920=0"

"D=8^2+4*1*1920=64+7680=7744"

"x1=(-8+88)\/2=40"

"x2=(-8-88)\/2=-48"

Choose the positive value, hence the speed of the passenger train was 40 km/h.